# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

### v3

@@ Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions, written in a minimalist Mathematical Programming Language (MPL), and systematically explores a configurable region of their deductive neighborhood. A defining feature of the architecture is its unified hash-based inference engine, which executes both algebraic manipulations and deterministic logical transformations. Definitions are compiled into a distributed grid of simple Logic Blocks (LBs) that exchange messages; whenever the premises of an inference rule unify, a new fact is emitted with full provenance to its sources, yielding replayable, auditable proof graphs. Experimental validation is performed on Elementary Number Theory (ENT) utilizing a batched execution strategy. Starting from foundational axioms and definitions, the system first develops first-order Peano arithmetic, which is subsequently applied to autonomously derive and prove Gauss’s summation formula as a main result. To manage combinatorial explosion, GL algorithmically enumerates conjectures and applies normalization, type constraints, and counterexample (CE) filtering. On commodity hardware, an end-to-end run completes in under 7 minutes. Generated proofs export as navigable HTML so that every inference step can be inspected independently. We outline a hardware-software co-design path toward massively parallel realizations and describe future integration with large language models (LLMs) for auto-formalization and conjecture seeding. The Python, C++, and MPL code to reproduce these experiments, along with the full proof graphs in HTML as well as machine-readable text format, are available in the project’s GitHub repository at github.com/Generative-Logic/GL commit 1771330 and are permanently archived at doi:10.5281/zenodo.17206386.
@@ 2 The GL Workflow
    Verifiable Output. Once the execution run is complete, GL generates a set of HTML files that serve as the final, verifiable output for visualization and debugging. These files contain a human-readable list of all proven theorems. Crucially, each step in every proof is hyperlinked to its justification—either an initial definition or a previously proven theorem—creating a fully auditable and transparent proof graph that can be independently verified.
@@
In the current implementation, simple facts such as addition and multiplication tables are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact — for example, that 3 × 4 = 12 — is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level arithmetic facts, effectively allowing the system to bootstrap its own fact base. However, GL currently lacks case differentiation — the ability to reason by cases such as trichotomy — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts.
@@ 3.4 CE Filter
Upcoming extension - Incubator.
In the current implementation, simple facts used by the CE filter are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery, regardless of the domain under exploration. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level facts, effectively allowing the system to bootstrap its own fact base. This tool becomes increasingly necessary as the complexity of the target domain grows: what counts as a "simple" fact in elementary arithmetic is a straightforward table entry, but in richer domains, the corresponding ground-level facts can themselves be non-trivial statements requiring structured derivation. Without automation, the manual curation of such facts quickly becomes a bottleneck. However, GL currently lacks case differentiation — the ability to reason by cases — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts, across any domain expressible in MPL.
@@ 3.7.2 The Interactive Proof Graph
Figure 6: A screenshot of a proof chapter (chapter81.html), showing several inference steps.
@@ 4 Case Study: Elementary Number Theory
To assess the GL system, we conducted a two-stage evaluation encompassing both Peano arithmetic and Gauss’s summation formula. The experiments were performed on standard consumer hardware (a Dell G16 7630 equipped with 32 logical cores and 32 GB of RAM). The complete execution finished in roughly 407 seconds, or just under seven minutes. A detailed breakdown of the time spent in each phase across the two batches is provided in Table 1. Notably, rather than the standard fractional form, GL autonomously derives the division-free variant of Gauss’s formula, expressing the product of $n$ and its successor as twice the sum:
@@
The conjecturer produces 406 candidates in 21 seconds. The CE filter processes all 406 conjectures in 14.6 seconds, passing 55 survivors to the prover. The prover runs 17 iterations (2 warm-up, 15 main) and finishes in 24.6 seconds, establishing 32 theorems. These cover the foundational arithmetic laws of Peano arithmetic, including commutativity and associativity of addition and multiplication, and distributivity. The 64 theorem entries written to file represent mirrored pairs — each theorem together with its symmetric counterpart. The total runtime of the native executable for this batch is 39.4 seconds.
@@
The Gauss batch begins by generating 363 conjectures in 106 seconds — the larger conjecture creation time reflecting the richer definition set, which now includes intervals, sequences, limit constructs, and the fold operator. Before the prover starts, the anchor connection mechanism automatically attaches the Gauss anchor to the Peano anchor, injecting the implication AnchorGauss $\Rightarrow$ AnchorPeano into the theorem list. This ensures that all 64 theorems proven in Batch 1 are available as ground knowledge for the Gauss prover.
@@
The CE filter runs for 192.5 seconds — the dominant cost in this batch — passing 29 conjectures to the prover. At the start of Phase 2, the prover broadcasts the 64 inherited Peano theorems across all 70 active memory blocks, making prior knowledge immediately available to every Logic Block. The prover then runs 30 iterations (2 warm-up, 28 main) and finishes in 47.1 seconds, proving 10 theorems including the flagship result: Gauss’s summation formula. The total runtime of the native executable for this batch is 239.7 seconds.
@@
Phase Peano (s) Share (%) Gauss (s) Share (%)
Conjecturer 21.3 5.2 106.0 26.0
CE filter 14.6 3.6 192.5 47.3
Prover 24.6 6.0 47.1 11.6
Total (executable) 39.4 9.7 239.7 58.8
Overall wall-clock: 407 seconds
@@
The runtime distribution shifts considerably between batches. In the Peano batch the three phases are roughly balanced; in the Gauss batch the CE filter becomes the dominant cost at 80% of the executable runtime, reflecting the increased complexity of the fact base required to screen a richer conjecture space. The prover, by contrast, remains tractable at under 50 seconds even with 64 inherited theorems broadcast into the grid. Conjecture creation accounts for the bulk of the Python-side runtime and remains the primary target for the planned migration to C++. These figures should be read as a snapshot of a initial prototype prototype; substantial headroom remains across all phases.
@@ 5.1 Contributions
Every reasoning step in GL is a symbolic manipulation based on a hash-table lookup. The final output is a complete, auditable proof graph where each inference is explicitly linked to a prior definition or a previously proven theorem. This provides a level of trust and verifiability that is not available in probabilistic systems. The correctness of a proof can be independently checked by a simple, external script, without needing to trust the internal machinery of GL itself.
@@ 6 Future Work, Roadmap and Potential Impact
With the Gauss summation formula proven and shipped, GL has completed its first two-batch ENT run. Peano delivered algebraic manipulation; Gauss delivered logical transformation. That closes one phase. The next milestone is the Fundamental Theorem of Arithmetic (FTA).
The road to FTA is not a continuation of the same work at larger scale. It requires three infrastructure extensions and one final major algorithmic extension, arriving in a specific order dictated by their dependencies.
6.1 Pruner
GL now carries a growing mass of proven theorems. That is good for expressive power but bad for focus. The pruner’s task is to select a minimal core subset from that mass such that all remaining theorems follow directly from it. This is not merely a runtime optimization. It is a structural compression step: it removes noise from theorem selection, improves proof trace readability, and ensures that the HTML proof graph reflects the real backbone of a result rather than drowning it in nearby consequences. As the theorem base grows, this compression step becomes essential for maintaining clarity and performance.
6.2 External Proof Checker
Internal proof generation is not sufficient once GL begins carrying heavier mathematical claims. An external checker creates a hard boundary between proving and verifying. GL generates proof objects; the checker validates them independently, without access to the internal machinery of the prover. This improves trust, makes debugging cleaner, and gives the project a stronger research posture. When the prover is modified, the checker remains the stable witness. When something fails, the boundary makes it immediately clear whether the issue lies in generation or in verification.
@@ 6.3 Case Differentiation
This is the last missing tool in GL’s mathematical stack, and arguably the most consequential. Peano gave GL algebraic manipulation. Gauss gave GL logical transformation. Case differentiation is the third pillar.
@@
6.4 Incubator
The Incubator becomes possible once case differentiation is available, because it depends on that tool to generate its outputs reliably. The incubator autogenerates ground-level fact tables — the machine-generated equivalent of multiplication tables — and feeds them into the CE filter as fuel. The CE filter grows stronger in proportion to the richness of its fact stock. The incubator is therefore a throughput engine: it manufactures the raw material that allows the CE filter to operate at higher quality and higher speed, without requiring manual curation of facts as the target domain grows.

### v4

@@ Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions written in a minimalist Mathematical Programming Language (MPL) and systematically explores a configurable region of their deductive neighborhood. Definitions are compiled into a distributed grid of Logic Blocks (LBs) that communicate via a unified hash-based inference engine; whenever the premises of a rule unify, a new fact is emitted with full provenance, yielding replayable, auditable proof graphs. The pipeline includes an Incubator that auto-generates ground-level fact tables, a Compressor that eliminates post-proof redundancy, and an independent external Verifier (34,320 checks, zero failures). Experimental validation on Elementary Number Theory develops Peano arithmetic from axioms and autonomously derives Gauss’s summation formula. On commodity hardware, the core proving pipeline completes in under one minute; the full run including Incubator fact generation finishes in approximately ten minutes. The Incubator output further reveals that GL can perform concrete numerical calculations — each result a proved theorem with full provenance — opening a path toward a full-provenance Computer Algebra System (CAS). Generated proofs export as navigable HTML for independent inspection. Code, proof graphs, and reproduction instructions are available at github.com/Generative-Logic/GL commit 6e5b9a4 and archived at doi:10.5281/zenodo.17206386.
@@ 2 The GL Workflow
Figure 1: GL pipeline architecture. MPL definitions feed both the Incubator (Step 0) and the main pipeline. Fact tables produced by the Incubator are consumed by the CE filter. The entire flow executes in ordered batches (e.g., IncubatorPeano $\to$ Peano $\to$ IncubatorGauss $\to$ Gauss).
  1. 0.
    Incubation (optional). Before the main pipeline runs, a dedicated tool called the Incubator auto-generates ground-level fact tables — such as addition and multiplication tables — directly from the axiomatic definitions. The Incubator uses GL’s own prover machinery with a contradiction-based proof strategy to conjecture and either prove or disprove ground-level arithmetic facts. Proved facts are fed into the CE filter of subsequent batches (Step 2), closing the bootstrap loop and making the pipeline self-sufficient from definitions down to ground facts. Each definition layer is incubated independently; previously established facts are reused without re-derivation.
@@
    Post-Proof Processing. After proof execution, two additional stages run. A Compressor performs post-proof redundancy elimination: it builds per-theorem proof graphs and applies greedy multi-pass elimination, retaining only the minimal set of essential theorems from which all others are derivable. The remaining theorems and their proof chains are then passed to the output stage.
  8. 7.
    Verifiable Output. Once the execution run is complete, GL generates a set of HTML files that serve as the final, verifiable output for visualization and debugging. These files contain a human-readable list of all proven theorems. Crucially, each step in every proof is hyperlinked to its justification—either an initial definition or a previously proven theorem—creating a fully auditable and transparent proof graph that can be independently verified. In addition, an independent external Verifier reads the processed proof graph and checks every inference step against 25 proof-tag-specific validators, providing a hard trust boundary between proof generation and proof verification.
@@ 3.4 CE Filter
Incubator — automated fact generation.
Ground-level facts used by the CE filter are generated automatically by a dedicated tool called the Incubator. Each such fact — for example, that $3+4=7$ or that the successor of $2$ is $3$ — is itself a simple theorem, and the Incubator uses GL’s own prover machinery to conjecture and either prove or disprove it. The Incubator is guided by dedicated configuration files (one per domain batch) and employs a contradiction-based proof strategy: for each conjecture, the negation of the head is attached to the proof chain, and a detected contradiction confirms the original statement. This tool becomes increasingly necessary as the complexity of the target domain grows: what counts as a “simple” fact in elementary arithmetic is a straightforward table entry, but in richer domains, the corresponding ground-level facts can themselves be non-trivial statements requiring structured derivation.
@@ 3.7.2 The Interactive Proof Graph
Figure 7: A screenshot of a proof chapter (chapter48.html), showing several inference steps.
@@ 4 Case Study: Elementary Number Theory
To assess the GL system, we conducted a four-batch evaluation encompassing Peano arithmetic, Gauss’s summation formula, and their corresponding Incubator stages. The experiments were performed on standard consumer hardware (a Dell G16 7630 equipped with 32 logical cores and 32 GB of RAM). The complete execution finished in roughly 602 seconds, or approximately ten minutes. The pipeline includes, beyond conjecture generation, CE filtering, and proof execution, an Incubator for auto-generating ground-level fact tables, a Compressor for post-proof redundancy elimination, and a Verifier for independent proof validation. A detailed breakdown of the time spent in each phase across all batches is provided in Table 1. Notably, rather than the standard fractional form, GL autonomously derives the division-free variant of Gauss’s formula, expressing the product of $n$ and its successor as twice the sum:
@@
The conjecturer produces 465 candidates in 3.2 seconds. The CE filter processes all 465 conjectures through two fact files in 11.0 seconds, passing 62 survivors to the prover. The prover runs 17 iterations (2 warm-up, 15 main) and finishes in 9.0 seconds. The Compressor then reduces the proved theorems to 27 essential (removing 56 redundant), covering the foundational arithmetic laws of Peano arithmetic, including commutativity and associativity of addition and multiplication, and distributivity. The total runtime of the native executable for this batch is 26.3 seconds.
@@
The Gauss batch begins by generating 363 conjectures in 3.6 seconds — the larger conjecture creation time reflecting the richer definition set, which now includes intervals, sequences, limit constructs, and the fold operator. Before the prover starts, the anchor connection mechanism automatically attaches the Gauss anchor to the Peano anchor, injecting the implication AnchorGauss $\Rightarrow$ AnchorPeano into the theorem list. This ensures that all 27 essential Peano theorems proven in Batch 1 are available as ground knowledge for the Gauss prover.
@@
The CE filter runs for 13.3 seconds, passing 72 conjectures to the prover. At the start of Phase 2, the prover broadcasts the 27 inherited Peano theorems across all active memory blocks, making prior knowledge immediately available to every Logic Block. The prover then runs 36 iterations (2 warm-up, 34 main) and finishes in 13.6 seconds, proving the flagship result: Gauss’s summation formula. The Compressor retains 39 globally essential theorems (removing 8 redundant). The total runtime of the native executable for this batch is 28.3 seconds.
Incubator
Before the main Peano and Gauss batches, the Incubator runs two dedicated batches — IncubatorPeano and IncubatorGauss — to auto-generate ground-level fact tables. The IncubatorPeano batch uses a 14-argument anchor (AnchorIncubator) with 9 scalar-typed arguments representing the elements $\{0,1,\ldots,8\}$ of a finite model. The Incubator employs a contradiction-based proof strategy: for each conjecture, the negation of the head is attached to the proof chain, and a detected contradiction confirms the original fact.
The IncubatorPeano batch processes 1575 conjectures and produces 1026 theorems in 323 seconds. The IncubatorGauss batch processes 243 conjectures and produces 177 theorems in 194 seconds, bringing the total incubator output to 1203 theorems. The positive facts cover successor, addition, and multiplication tables (e.g., $s(0)=1$, $0+5=5$, $2\times 3=6$); the contradictions establish negative facts (e.g., $s(0)\neq 4$, $0\neq 1$). Unresolved conjectures are negative facts whose proof requires intermediate values exceeding the model range — for instance, proving $5+4\neq 10$ requires computing $5+4=9$, which falls outside the $\{0,\ldots,8\}$ model. Enlarging the anchor with additional elements would resolve these at the cost of increased combinatorial load. The proved facts are fed into the CE filter of subsequent main batches, closing the bootstrap loop from definitions to ground facts.
The Incubator is heuristic in nature. Its primary tool is back-reformulation: a proved statement of the form $2+2=x\implies x=4$ is reformulated into the direct operator fact $2+2=4$. This step is not a valid mathematical conclusion in the strict sense, as it omits the existence proof that the operator application $2+2$ yields a unique value — a gap that becomes particularly relevant for compound operators such as fold. The real risk is therefore not that an incorrect fact enters the proof graph, but that a correct conjecture is wrongly refused by the CE filter on the basis of an erroneous Incubator fact. Users must ensure that existence is guaranteed in pre-analysis — a due-diligence step. Consequently, the proof graph may be incomplete: some valid theorems could be missing because their conjectures were filtered prematurely. However, whatever is inside the proof graph is still correct — Incubator facts never enter the main proof path, and the prover itself remains sound.
Compressor
After proof execution, the Compressor identifies a minimal essential subset of theorems. In Phase 1, it creates an independent Logic Block for each of the proved theorems, loads all theorems as hash-memory inference rules, and runs hash bursts to extract a per-theorem proof graph recording which other theorems were used in each derivation. In Phase 2, it performs greedy multi-pass elimination: theorems are sorted by usage frequency (least-used first) and tentatively removed; a forward-reachability check (isDerivable) verifies that all surviving theorems remain derivable from the reduced set. The process iterates until a fixed point is reached. For the ENT case study, the Peano Compressor reduces 83 theorems to 27 essential (removing 56 redundant), and the Gauss Compressor retains 39 globally essential theorems (removing 8) — without loss of deductive coverage.
Verifier
An external, independent Verifier reads the processed proof graph and validates every inference step. The Verifier is implemented in pure Python with zero coupling to the C++ prover codebase: it shares no code, no data structures, and no runtime dependencies with the proving engine. It checks each proof line against 25 tag-specific validators covering the full range of proof mechanisms: implication, expansion, disintegration, equality (substitution, transitivity, symmetry), recursion (induction step), theorem application, mirrored and reformulated forms, integration (reformulation and expansion), anchor handling, contradiction, and several special-purpose tags. For the ENT case study, the Verifier performs 34,320 checks across all proof chapters (32,668 incubator, 1,652 main) and reports zero failures.
@@
Phase Inc. Peano (s) Peano (s) Inc. Gauss (s) Gauss (s)
Conjecturer 0.2 3.2 0.4 3.6
CE filter 11.0 13.3
Prover 321.2 9.0 171.7 13.6
Compressor incl. incl.
Total (executable) 323.1 26.3 194.5 28.3
Verifier: 34,320 checks (32,668 + 1,652), 0 failures
Overall wall-clock: $\sim$602 seconds
@@
The runtime distribution separates cleanly into two regimes. The core proving pipeline (Peano + Gauss) completes in approximately 55 seconds of executable time, with conjecture creation in single-digit seconds (C++ conjecturer), CE filtering in 11–13 seconds per batch, and proof execution in 9–14 seconds. The Incubator batches account for the majority of the overall wall-clock time (517 of 602 seconds), reflecting the cost of generating ground-level fact tables from first principles across a nine-element model. Int-path data structures replace string-based operations in the prover’s hot path, and a static memory allocation pipeline serves the hash engine. The pipeline is self-bootstrapping: the Incubator auto-generates the 1203 ground-level facts that the CE filter consumes, the Compressor reduces each batch’s theorem mass to a minimal essential core (27 Peano, 39 globally after Gauss), and the Verifier confirms zero failures across all 34,320 checks.
The Incubator’s dominance of the overall runtime is less problematic than it appears. Ground-level fact tables do not need to be regenerated on every run: once the multiplication table is proved, it does not need to be reproved when the system moves on to algebraic geometry. In practice, previously incubated facts can be supplied externally, and only facts for newly introduced definitions require fresh incubation. The IncubatorGauss batch, which covers the fold and interval definitions added beyond Peano, processes only 243 conjectures in 194 seconds — compared to 517 seconds for both Incubator batches combined — an order of magnitude fewer than the 1575 of IncubatorPeano, which carries the full cost of bootstrapping basic arithmetic from scratch. As the system grows, each new definition layer will incubate only its own ground-level facts on top of the existing base. The Incubator will likely remain the most expensive pipeline stage, but its cost is incremental rather than cumulative, and it can be amortized across runs.
Beyond its role in the pipeline, the Incubator output reveals an unexpected capability: GL is not only proving properties but computing concrete numerical results, each carrying full provenance back to the axioms. This observation motivates the following section.
5 From Proofs to Calculations
The Incubator results reveal a capability that was not an explicit design goal but emerges naturally from GL’s architecture: the system can perform concrete numerical calculations. When GL proves that $2+3=5$ or that $\text{fold}(+,[0\ldots 4])=\sum_{i=0}^{4}i=10$, it is not merely verifying a pre-stated identity — it is computing the result, deriving a specific numerical value from the axiomatic definitions of the operators involved. Every intermediate step of that computation is recorded in the proof graph with full provenance to its source axioms.
Three levels of calculation capability are demonstrated in the current Incubator output:
  1. 1.
    Evaluation of operators. GL evaluates successor, addition, and multiplication on concrete arguments — e.g., $s(2)=3$, $4+1=5$, $2\times 3=6$ — by chasing the recursive definitions to their base cases and assembling the result. Each such fact is a proved theorem.
  2. 2.
    Application of formulas. Operator results compose: the system derives facts such as $s(a+b)=c$ by combining the definitions of successor and addition, effectively applying a formula built from multiple operators to concrete values.
  3. 3.
    Execution of multi-stage algorithms. The fold operator encapsulates an iterative accumulation over an interval — the logical equivalent of a for loop. When GL evaluates a fold instance, it executes a multi-step algorithmic process: establishing the interval bounds, iterating the accumulator through each element, and arriving at the final value. The proof graph records every iteration, producing a fully auditable trace of the computation.
A defining feature of this capability is that no separate calculation mode exists. The same MPL definitions, the same hash-based inference engine, and the same Logic Block grid that prove abstract properties such as the commutativity of addition also compute that $2+3=5$. A property is verified and a calculation is executed by the same mechanism: proof is calculation, and calculation is proof. This unification means that every computed value carries the same provenance guarantees as every proven theorem — a chain of justified inference steps leading back to the foundational axioms.
It is important to distinguish the Incubator’s current heuristic shortcut from the intended calculation path. The Incubator operates before the main proof campaign: the theorems it would need for a fully rigorous derivation are not yet conjectured, let alone proved, so it resorts to back-reformulation as a pragmatic workaround (Section 4). A future calculator, by contrast, operates after the full theorem base has been established. It consumes the complete, verified output of GL and executes calculations using only proved inference rules — no heuristic reformulation step is involved. The calculation path is therefore as rigorous as the proof path itself.
This unification opens a path toward what might be called a full-provenance symbolic procedural Computer Algebra System (CAS): a system in which every computed result is a theorem and every theorem can be a computed result, with each intermediate step traceable to the foundational axioms. The current implementation is modest — single-digit arithmetic on a nine-element model, with the Incubator requiring several minutes to generate its fact tables — but the architecture does not impose any ceiling on the complexity of the computations it can trace. What is demonstrated here is a first, very modest, but nonetheless real appearance of calculation capability within a fully auditable reasoning architecture.
@@ 6.1 Contributions
Every reasoning step in GL is a symbolic manipulation based on a hash-table lookup. The final output is a complete, auditable proof graph where each inference is explicitly linked to a prior definition or a previously proven theorem. This provides a level of trust and verifiability that is not available in probabilistic systems. The correctness of a proof can be independently checked by a simple, external script, without needing to trust the internal machinery of GL itself. This principle is now realized concretely through an external Verifier that reads only the processed proof graph and checks every inference step against 25 tag-specific validators, achieving zero failures across all checks in the ENT case study.
@@ 7 Future Work, Roadmap and Potential Impact
The next milestone is the Fundamental Theorem of Arithmetic (FTA). The road to FTA requires one major algorithmic extension — case differentiation — and a sustained engineering campaign.
@@ 7.1 Case Differentiation
Case differentiation is the next missing tool in GL’s mathematical stack. Peano gave GL algebraic manipulation. Gauss gave GL logical transformation. Case differentiation is the next pillar.
@@
The mechanism fits naturally into GL’s existing infrastructure. When the prover encounters a branching point, it emits sub-facts tagged with an explicit or-tag that identifies the branch they belong to. Each branch proceeds independently inside its own Logic Blocks, and the results carry their or-tags through every subsequent inference step. When all branches of an OR event deliver a common conclusion, the or-tags are stripped and the conclusion is promoted to an unconditional fact.
The conjecture generator contributes the second half. For every proved implication $\forall x.\;A(x)\!\Rightarrow\!\lnot B(x)$, a negation copy $\forall x.\;B(x)\!\Rightarrow\!\lnot A(x)$ is automatically generated and submitted for proof. If both an implication and its reverse are established, the system recognizes $\lnot(A(x)\land B(x))$ — equivalently $\lnot A(x)\lor\lnot B(x)$ — and broadcasts it as an OR theorem, making it available as a branching point for subsequent proofs.
As an illustration, consider Euclid’s lemma: if a prime $p$ divides a product $ab$, then $p\mid a$ or $p\mid b$. GL would first prove the two complementary implications — roughly, $p\mid a$ implies $p\nmid b$ is not required” and vice versa — then detect the complementary pair and emit the disjunction. Any later proof that requires case analysis on divisibility by a prime can then branch on this OR theorem, pursue each branch independently, and merge the results once both branches converge.